OpenStudy (anonymous):

Karl is putting chairs into rows. If he puts 7 chairs in each row he will make 3 rows and have 3 chairs left over. If he instead puts 8 chairs in each row, how many rows of chairs can he make?

OpenStudy (anonymous):

No. of chairs is 7x3+3=24, so if he puts the chairs into rows of 8, he will get 24/8=3 rows.

OpenStudy (asadkarim7):

7 x3 =21 the chairs left = 3 21+3 = 24 24 / 8 3 rows

OpenStudy (anonymous):

correct